OpenStudy (anonymous):

Lesson 10: Radical Functions and Rational Exponents. Has anyone taken it? Please I need help or I'm gonna fail the class :(

OpenStudy (amoodarya):

whats the question ?

OpenStudy (anonymous):

There is more than a few... 1.Which of the following is equal to sqrt18x^9y^4? 2. What is the simplest form of the expression sqrt2-sqrt6 over sqrt2+sqrt6

OpenStudy (amoodarya):

\[\sqrt{18x^{9}y^{4}}=\sqrt{9 \times 2 x^{9}y^{4}}=3\times y^2 \sqrt{2 x^9} =3y^2x^4\sqrt{2x}\]

OpenStudy (amoodarya):

\[\frac{ (\sqrt{2}-\sqrt{6}) }{ (\sqrt{2}+\sqrt{6}) } =\] \[\frac{ (\sqrt{2}-\sqrt{6}) }{ (\sqrt{2}+\sqrt{6}) } \times \frac{ (\sqrt{2}-\sqrt{6}) }{ (\sqrt{2}-\sqrt{6}) } =\] \[=\frac{ (\sqrt{2}-\sqrt{6})^2 }{ (2-6) } =\frac{ (2+6-2\sqrt{6}) }{ (2-6) } =\frac{ (8-2\sqrt{6}) }{ (-4) }\] simplify

OpenStudy (anonymous):

how do I simplify that?